X Company uses an activity-based costing overhead allocation system. It has identified three activities - Setup, Purchasing, Packaging - and their respective cost drivers - setup hours, purchase orders, packages. Budgeted activity costs in 2018 were as follows:

Activity Budgeted Cost  
Setup $104,200      
Purchasing 55,400      
Packaging 60,500      

The following 2018 cost driver information is available for one of its products, Product X, and for all of its products (including Product X):

Product X All Products
setup hours 31,900     744,286    
purchase orders 51,900     1,385,000    
packages 66,600     864,286    


How much overhead was allocated to Product X in 2018 [round overhead rate(s) to two decimal places]?
Answer #1

Overhead rate=Respective cost/Respective activity

Overhead rate for:

Setup=(104200/744286)=$0.14 per setup hour

Purchasing=(55400/1,385,000)=$0.04 per purchase order

Packaging=(60500/864,286)=$0.07 per package

Hence overhead allocated to Product X=(0.14*31900)+(0.04*51900)+(66,600*0.07)

=$11204.